OpenStudy (anonymous):

simplify the following expression y^2-4y-21/y-7

OpenStudy (nurali):

\[\frac{ y^2-4y-21 }{ y-7 }\] \[\frac{ y^2-7y+3y-21 }{ y-7 }\] \[\frac{ (y-7)(y+3) }{ (y-7) }\] \[(y+3)\]